# GoldCoin: Grounding Large Language Models in Privacy Laws via Contextual Integrity Theory

### v1

@@ Abstract
Privacy issues arise prominently during the inappropriate transmission of information between entities. Existing research primarily studies privacy by exploring various privacy attacks, defenses, and evaluations within narrowly predefined patterns, while neglecting that privacy is not an isolated, context-free concept limited to traditionally sensitive data (e.g., social security numbers), but intertwined with intricate social contexts that complicate the identification and analysis of potential privacy violations. The advent of Large Language Models (LLMs) offers unprecedented opportunities for incorporating the nuanced scenarios outlined in privacy laws to tackle these complex privacy issues. However, the scarcity of open-source relevant case studies restricts the efficiency of LLMs in aligning with specific legal statutes. To address this challenge, we introduce a novel framework, GoldCoin, designed to efficiently ground LLMs in privacy laws for judicial assessing privacy violations. Our framework leverages the theory of contextual integrity as a bridge, creating numerous synthetic scenarios grounded in relevant privacy statutes (e.g., HIPAA), to assist LLMs in comprehending the complex contexts for identifying privacy risks in the real world. Extensive experimental results demonstrate that GoldCoin markedly enhances LLMs’ capabilities in recognizing privacy risks across real court cases, surpassing the baselines on different judicial tasks.
@@
Wei Fanthanks:  Equal Contribution, Haoran Li, Zheye Deng, Weiqi Wang, Yangqiu Song Department of Computer Science and Engineering, HKUST, Hong Kong SAR, China {wfanag, hlibt, zdengah}@connect.ust.hk, {wwangbw, yqsong}@cse.ust.hk
@@ 1 Introduction
Intuitively, we consider applying the wealth of real-world scenarios contained in legal statutes and case law to address the limitation. However, converting legislation into an actionable framework remains a significant challenge. Previous efforts have involved translating legislation into logical languages Lam et al. (2009); DeYoung et al. (2010); Robaldo et al. (2020), yet this method heavily relies on expert annotation and struggles to adapt to legislative changes or scale across different privacy laws. The recent emergence of LLMs OpenAI (2022); Touvron et al. (2023b); Anthropic (2024), has introduced new potential for addressing the problem. Specifically, legal LLMs like LawGPT Zhou et al. (2024), Lawyer LLaMA Huang et al. (2023), ChatLaw Cui et al. (2023) have all leveraged the vast existing statutes and cases to assist public in general legal tasks.
@@ 2.3 LLMs for Instruction Generation
A series of works have explored the use of LLMs for data generation Meng et al. (2023); Liu et al. (2022); Wang et al. (2021); Schick and Schütze (2021). Recent studies have particularly concentrated on enhancing instruction generation Honovich et al. (2022a); Zhou et al. (2023); Singh et al. (2023); Honovich et al. (2022b); Wang et al. (2023) to improve zero-shot Ye et al. (2022) and few-shot Brown et al. (2020); Wei et al. (2021) learning, abstraction reasoning Wang et al. (2024) capabilities, as well as the instruction-following proficiency of LLMs. Inspired by it, our approach utilizes the strong generative capabilities of GPT-4 to address the case scarcity based on contextual integrity theory by instructing it to generate datasets and align LLMs with privacy laws for judicial.
@@ 3 Method
Legal judgments on privacy violations typically involve two tasks Lam et al. (2009): (1) Applicability, assessing whether the privacy law $\mathcal{L}$ applies to the case background $s$; and (2) Compliance, determining if the transmission described in $s$ compliant with $\mathcal{L}$. In this section, we introduce GoldCoin, which applies contextual integrity theory to generate synthetic cases. After postprocessing the instances, we instruction-tune LLMs and evaluate their performance of the above two tasks. The overview of our pipeline is shown in Figure 3.
@@ 3.1 Legal Statute Preprocessing
To evaluate the effectiveness of GoldCoin, we apply it to the U.S. Health Insurance Portability and Accountability Act (HIPAA) Privacy Rule. Initially, we dump the content of the HIPAA Privacy Rule from the official Code of Federal Regulations (CFR) website [^2] https://www.ecfr.gov/current/title-45/subtitle-A/subchapter-C. We then transform the textual data into a structured graph $\mathcal{G}$, comprising nodes $\mathcal{V}$ that represent sections and two types of relations $\mathcal{E}$. These relationships are identified as subsume, denoting hierarchical relationships (e.g., (164.502(a), subsume, 164.502)), and refer, indicating cross-references between sections (e.g., (164.502(a)(1)(ii), refer, 164.504(b))). Each node consistes of a labeled identifier and the paragraph content. We start from each leaf node $v^{l}_{i}$ and recursively identify all parent nodes $\{v^{l-1}_{i},v^{l-2}_{i},\ldots,v^{0}_{i}\}$ where $v^{0}_{i}$ is the root node. Then we aggregate their content, as depicted in Figure 2 and refer to each such path as a norm. These norms can be categorized into three types: permit $n^{+}$, forbid $n^{-}$, and others. The first two categories describe the permissions and prohibitions regarding information transmission under the law, while the last category contains general definitions, exceptions, and requirements. We leverage GPT-4 to classify and label these leaf norms and filter the permit and forbid norm for the subsequent generation steps. The examples of permit and forbid norms are shown in the upper part of Figure 8 and Figure 9, respectively. All the details of the HIPAA Privacy Rule and preprocessing are depicted in Appendix B.
@@ 3.2 Law-Grounded Case Generation
Given the seed norm $n_{i}$ and the conclusion $c_{i}$ which correspond to the norm type (i.e., permit, forbid), we manually build the instructions combined with $n_{i}$ for background generation. To ensure the generation of background narratives that align with $n_{i}$ and preserve the integrity of the privacy information transmission context, we construct a detailed prompt (see Section C.1) that includes the description of the key features in contextual integrity, such as entities, roles, information type, and transmission principles.
@@ 3.3 Case Postprocessing
To mitigate the reduction in semantic complexity and alignment robustness caused by similar features across different case backgrounds, we implement the methodology from prior studies Wang et al. (2023, 2024) to promote background diversity. We calculate the ROUGE-L Lin (2004) similarity for each case background against others, ranking them accordingly. For each norm, we select the case with the highest ROUGE-L score to ensure optimal diversity.
@@ 3.4 Real Court Case Collection
To rigorously validate the grounding efficacy of GoldCoin, we retrieve all relevant real court cases featuring the “HIPAA Privacy Rule" from the Caselaw Access Project (CAP) [^3] https://case.law/, developed by Harvard Law School. These cases are systematically processed and distilled into the same format as synthetic cases, utilizing both GPT-4 and manual annotation. Additionally, as delineated in Section 3.2, our methodology focuses exclusively on constructing cases applicable to the HIPAA Privacy Rule, because privacy cases unrelated to HIPAA are readily available. To provide a negative training and testing set for the applicability task, we also curate a collection of cases that, while closely related to privacy violations, do not apply to HIPAA. The details of CAP are explained in Appendix D.
@@ 3.5 Instruction and Response Compilation
We construct responses for applicability task following two steps: first, extracting context-specific features; second, determining HIPAA applicability:
@@ 4.2 CAP Real Court Cases
Due to reasons outlined in Section 3.4, we directly collect cases irrelevant to HIPAA as negative training examples. We select cases tagged with “Privacy Violation” using the “Most Relevant First” search function on the CAP website, gathering 309 non-applicable cases for training. For evaluation, after a combined screening by GPT-4 and human experts, we identify 107 real court cases relevant to HIPAA, serving as the ground truth for the compliance task. Correspondingly, we also sample an equivalent number of HIPAA-irrelevant cases and combine them with the 107 cases to form the test set for the applicability task. Ultimately, we combine synthetic and real cases to create the GoldCoin-hipaa dataset.
@@ 5.2 Overall Performance
We first analyze the performance of four LLMs in determining the HIPAA applicability of real court cases sourced from CAP. Our results demonstrate that GoldCoin can align the LLMs with the comprehensive understanding of the HIPAA Privacy Rule, exceeding all baseline methods. Notably, MPT-7B, which performed near-random levels (Acc  50%, Ma-F1  50%), see substantial improvements with our method—accuracy and Macro F1-scores increase by 12.62% and 11.81%, respectively, compared to the zero-shot setting. Meanwhile, Mistral-7B and Llama2-13B tuned with our framework, achieve exceptional accuracy rates of 97.66% and 99.53%, respectively, even attaining 100% in “Not applicable” category (see Table 12), surpassing the performance of ChatGPT and GPT-4. We observe that MPT-7B, when trained exclusively with “Direct Prompt,” exhibits only a limited improvement of 2.49% in Ma-F1. This underscores the integration of contextual features is crucial for decomposing and deeply understanding legal case topics. Additionally, our results indicate that merely continuing to train LLMs on legal statutes results in limited effectiveness and even leads to diminished performance in determining applicability (e.g., MPT-7B $\downarrow$10.8%).
@@
Our GoldCoin framework introduces multi-step simulated trial instructions, effectively aligning LLMs with privacy law and enhancing their reasoning capabilities on compliance tasks. It significantly improved Macro-F1 scores across several models: MPT-7B (17.87%), Llama2-7B (12.45%), Mistral-7B (17.96%), and Llama2-13B (8.12%) compared to the zero-shot setting. Mistral-7B, specifically tuned on our dataset, excels in precision for both “permit” and “forbid” cases, surpassing ChatGPT and approaching GPT-4’s performance. However, using “Direct Prompt” result in a notable decline for Mistral-7B, from 66.98% to 51.75%, indicating limited grounding ability. Direct training on abstract legal concepts leads to reasoning confusion, as seen with Llama2-7B, which tends to misclassify cases as “forbid,” (see Table 13). Our results reaffirm the high quality of cases generated under contextual integrity theory and the feasibility of the reasoning pipeline for adjudicating privacy law cases.
@@ 5.3 Ablation Study
We conduct ablation studies to assess the effect of contextual feature filters. After generating case backgrounds, we retain all cases including those that lacked key features (e.g., sender, recipient) of contextual integrity. The results, denoted as ($\diamond$ w/o Feature F), reveal significant performance declines. Specifically, there is a drop of 3.26% and 2.36% in the applicability and compliance tasks, respectively, for Llama2-13B (see Table 3). These findings demonstrate the importance of feature integrity.
@@
We remove the diversity ranking ($\diamond$ w/o Diversity R) and randomly sample cases for each norm. Low diversity often results in high similarity among cases, such as in the roles of entities or specific categories of information. The lack of diversity can decrease the robustness of training, as demonstrated in Wang et al. (2024, 2023). This impact is further reflected in a 3.86% decline in the Macro F1-score for applicability judgments in Llama2-13B. Furthermore, we deactivate all of the above filters and ranking mechanisms ($\diamond$ w/o All Parts) and observe significant decreases across all language models, with Mistral-7B experiencing drops of 5.89% and 5.05% in each task respectively. These findings underscore the importance of enhancing the integrity, consistency, and diversity of generated cases.
@@ 6 Conclusion
Although our approach considers all permit and forbid norms within HIPAA, it does not account for interconnections between these norms. In practice, legal norms often contain cross-references and a single case may be adjudicated based on multiple norms Lam et al. (2009). Future work should construct cases based on multiple norms to better reflect real-world scenarios and potentially yield improvements. Additionally, we do not consider the few-shot setting due to multiple examples often exceeding the maximum input length of LLMs. For the selection of laws, we conduct experiments on HIPAA due to its prominence in the privacy domain and the relatively abundant availability of open-source cases, which can serve as ground truth for testing. We invite legal professionals with access to cases related to other privacy laws to contact us, as this would facilitate the extension of our approach to additional privacy regulations such as COPPA Aftab and Savitt (1999), GDPR Voigt and Von dem Bussche (2017), etc. Moreover, this paper primarily focuses on case generation, and we do not employ techniques such as retrieval-augmented generation Gao et al. (2024); Lewis et al. (2020) or vector embedding Douze et al. (2024) for retrieving relevant norms. We believe that dynamically indexing Liu (2022) and retrieving related norms, based on the statute graph constructed in Section 3.1, is a promising direction.
@@
We use the API provided by the Code of Federal Regulations official website to access the HIPAA Privacy Rule. We follow contextual integrity theory Nissenbaum (2004) to generate synthetic cases for constructing GoldCoin-hipaa, and manually remove cases that could potentially leak real-world private information. We follow the official usage and access rules of the Caselaw Access Project [^4] https://case.law/about/#usage-access during downloading relevant cases. Human evaluations and annotations are performed by two legal experts to review the quality of synthetic cases and remove cases that contain explicit content such as gore or violence. Annotations are compensated at 15 USD per hour, above the local minimum wage. To the best of our knowledge, this work complies with open-source agreements and does not pose risks of information leakage or other hazards.

### v2

@@ Abstract
Privacy issues arise prominently during the inappropriate transmission of information between entities. Existing research primarily studies privacy by exploring various privacy attacks, defenses, and evaluations within narrowly predefined patterns, while neglecting that privacy is not an isolated, context-free concept limited to traditionally sensitive data (e.g., social security numbers), but intertwined with intricate social contexts that complicate the identification and analysis of potential privacy violations. The advent of Large Language Models (LLMs) offers unprecedented opportunities for incorporating the nuanced scenarios outlined in privacy laws to tackle these complex privacy issues. However, the scarcity of open-source relevant case studies restricts the efficiency of LLMs in aligning with specific legal statutes. To address this challenge, we introduce a novel framework, GoldCoin  [^1] The code and data are available at https://github.com/HKUST-KnowComp/GoldCoin, designed to efficiently ground LLMs in privacy laws for judicial assessing privacy violations. Our framework leverages the theory of contextual integrity as a bridge, creating numerous synthetic scenarios grounded in relevant privacy statutes (e.g., HIPAA), to assist LLMs in comprehending the complex contexts for identifying privacy risks in the real world. Extensive experimental results demonstrate that GoldCoin markedly enhances LLMs’ capabilities in recognizing privacy risks across real court cases, surpassing the baselines on different judicial tasks.
@@
Wei Fanthanks:  Equal Contribution, Haoran Lithanks:  Corresponding Author, Zheye Deng, Weiqi Wang, Yangqiu Song Department of Computer Science and Engineering, HKUST, Hong Kong SAR, China {wfanag, hlibt, zdengah}@connect.ust.hk, {wwangbw, yqsong}@cse.ust.hk
@@ 1 Introduction
Intuitively, we consider applying the wealth of real-world scenarios contained in legal statutes and case law to address the limitation. However, converting legislation into an actionable framework remains a significant challenge. Previous efforts have involved translating legislation into logical languages Lam et al. (2009); DeYoung et al. (2010); Robaldo et al. (2020), yet this method heavily relies on expert annotation and struggles to adapt to legislative changes or scale across different privacy laws. The recent emergence of LLMs OpenAI (2022); Touvron et al. (2023b); Anthropic (2024) has introduced new potential for addressing the problem. Specifically, legal LLMs like LawGPT Zhou et al. (2024), Lawyer LLaMA Huang et al. (2023), ChatLaw Cui et al. (2023) have all leveraged the vast existing statutes and cases to assist public in general legal tasks.
@@ 2.3 LLMs for Instruction Generation
A series of works have explored the use of LLMs for data generation Meng et al. (2023); Liu et al. (2022); Schick and Schütze (2021); Wang et al. (2024a); Wang and Song (2024). Recent studies have particularly concentrated on enhancing instruction generation Honovich et al. (2022a); Zhou et al. (2023); Singh et al. (2023); Honovich et al. (2022b); Wang et al. (2023) to improve zero-shot Ye et al. (2022) and few-shot Brown et al. (2020); Wei et al. (2021) learning, abstraction reasoning Wang et al. (2024b) capabilities, as well as the instruction-following proficiency of LLMs. Inspired by it, our approach utilizes the strong generative capabilities of GPT-4 to address the case scarcity based on contextual integrity theory by instructing it to generate datasets and align LLMs with privacy laws for judicial.
@@ 3 Method
Legal judgments on privacy violations typically involve two tasks Lam et al. (2009): (1) Applicability, assessing whether the privacy law $\mathcal{L}$ applies to the case background $s$; and (2) Compliance, determining if the transmission described in $s$ compliant with $\mathcal{L}$. In this section, we introduce GoldCoin, which applies contextual integrity theory to generate synthetic cases. After postprocessing the instances, we instruction-tune LLMs and evaluate their performance in the above two tasks. The overview of our pipeline is shown in Figure 3.
@@ 3.1 Legal Statute Preprocessing
To evaluate the effectiveness of GoldCoin, we apply it to the U.S. Health Insurance Portability and Accountability Act (HIPAA) Privacy Rule. Initially, we dump the content of the HIPAA Privacy Rule from the official Code of Federal Regulations (CFR) website [^3] https://www.ecfr.gov/current/title-45/subtitle-A/subchapter-C. We then transform the textual data into a structured graph $\mathcal{G}$, comprising nodes $\mathcal{V}$ that represent sections and two types of relations $\mathcal{E}$. These relationships are identified as subsume, denoting hierarchical relationships (e.g., (164.502(a), subsume, 164.502)), and refer, indicating cross-references between sections (e.g., (164.502(a)(1)(ii), refer, 164.504(b))). Each node consists of a labeled identifier and the paragraph content. We start from each leaf node $v^{l}_{i}$ and recursively identify all parent nodes $\{v^{l-1}_{i},v^{l-2}_{i},\ldots,v^{0}_{i}\}$ where $v^{0}_{i}$ is the root node. Then, we aggregate their content, as depicted in Figure 2, and refer to each such path as a norm. These norms can be categorized into three types: permit $n^{+}$, forbid $n^{-}$, and others. The first two categories describe the permissions and prohibitions regarding information transmission under the law, while the last category contains general definitions, exceptions, and requirements. We leverage GPT-4 to classify and label these leaf norms and filter the permit and forbid norm for the subsequent generation steps. The examples of permit and forbid norms are shown in the upper part of Figure 8 and Figure 9, respectively. All the details of the HIPAA Privacy Rule and preprocessing are depicted in Appendix B.
@@ 3.2 Law-Grounded Case Generation
Given the seed norm $n_{i}$ and the conclusion $c_{i}$, which correspond to the norm type (i.e., permit, forbid), we manually build the instructions combined with $n_{i}$ for background generation. To ensure the generation of background narratives that align with $n_{i}$ and preserve the integrity of the privacy information transmission context, we construct a detailed prompt (see Section C.1) that includes the description of the key features in contextual integrity, such as entities, roles, information type, and transmission principles.
@@ 3.3 Case Postprocessing
To mitigate the reduction in semantic complexity and alignment robustness caused by similar features across different case backgrounds, we implement the methodology from prior studies Wang et al. (2023, 2024b) to promote background diversity. We calculate the ROUGE-L Lin (2004) similarity for each case background against others, ranking them accordingly. For each norm, we select the case with the highest ROUGE-L score to ensure optimal diversity.
@@ 3.4 Real Court Case Collection
To rigorously validate the grounding efficacy of GoldCoin, we retrieve all relevant real court cases featuring the “HIPAA Privacy Rule" from the Caselaw Access Project (CAP) [^4] https://case.law/, developed by Harvard Law School. These cases are systematically processed and distilled into the same format as synthetic cases, utilizing both GPT-4 and manual annotation. Additionally, as delineated in Section 3.2, our methodology focuses exclusively on constructing cases applicable to the HIPAA Privacy Rule because privacy cases unrelated to HIPAA are readily available. To provide a negative training and testing set for the applicability task, we also curate a collection of cases that, while closely related to privacy violations, do not apply to HIPAA. The details of CAP are explained in Appendix D.
@@ 3.5 Instruction and Response Compilation
We construct responses for the applicability task following two steps: first, extracting context-specific features; second, determining HIPAA applicability:
@@ 4.2 CAP Real Court Cases
Due to reasons outlined in Section 3.4, we directly collect cases irrelevant to HIPAA as negative training examples. We select cases tagged with “Privacy Violation” using the “Most Relevant First” search function on the CAP website, gathering 309 non-applicable cases for training. For evaluation, after a combined screening by GPT-4 and human experts, we identify 107 real court cases relevant to HIPAA, serving as the ground truth for the compliance task. Correspondingly, we also sample an equivalent number of HIPAA-irrelevant cases and combine them with the 107 cases to form the test set for the applicability task. Ultimately, we combine synthetic and real cases to create the GoldCoin-hipaa dataset. Appendix D details the collection and post-processing of CAP cases.
@@ 5.2 Overall Performance
We first analyze the performance of four LLMs in determining the HIPAA applicability of real court cases sourced from CAP. Our results demonstrate that GoldCoin can align the LLMs with the comprehensive understanding of the HIPAA Privacy Rule, exceeding all baseline methods. Notably, MPT-7B, which performed near-random levels (Acc  50%, Ma-F1  50%), see substantial improvements with our method—accuracy and Macro F1-scores increase by 12.62% and 11.81%, respectively, compared to the zero-shot setting. Meanwhile, Mistral-7B and Llama2-13B, tuned with our framework, achieve exceptional accuracy rates of 97.66% and 99.53%, respectively, even attaining 100% in the “Not applicable” category (see Table 13), surpassing the performance of ChatGPT and GPT-4. We observe that MPT-7B, when trained exclusively with “Direct Prompt,” exhibits only a limited improvement of 2.49% in Ma-F1. This underscores the integration of contextual features, which is crucial for decomposing and deeply understanding legal case topics. Additionally, our results indicate that merely continuing to train LLMs on legal statutes results in limited effectiveness and even leads to diminished performance in determining applicability (e.g., MPT-7B $\downarrow$10.8%).
@@
Our GoldCoin framework introduces multi-step simulated trial instructions, effectively aligning LLMs with privacy law and enhancing their reasoning capabilities on compliance tasks. It significantly improved Macro F1-scores across several models: MPT-7B (17.87%), Llama2-7B (12.45%), Mistral-7B (17.96%), and Llama2-13B (8.12%) compared to the zero-shot setting. Mistral-7B, specifically tuned on our dataset, excels in precision for both “permit” and “forbid” cases, surpassing ChatGPT and approaching GPT-4’s performance. However, using “Direct Prompt” results in a notable decline for Mistral-7B, from 66.98% to 51.75%, indicating limited grounding ability. Direct training on abstract legal concepts leads to reasoning confusion, as seen with Llama2-7B, which tends to misclassify cases as “forbid” (see Table 14). Our results reaffirm the high quality of cases generated under contextual integrity theory and the feasibility of the reasoning pipeline for adjudicating privacy law cases. Furthermore, to demonstrate that sample imbalance does not affect overall results, we apply oversampling to the forbid cases and assess the performance impact on the Mistral-7B and Llama2-13B models. Our findings indicate no significant change in the final performance, as detailed in Table 12 in Section F.4.
@@ 5.3 Ablation Study
We conduct ablation studies to assess the effect of contextual feature filters. After generating case backgrounds, we retain all cases, including those that lacked key features (e.g., sender, recipient) of contextual integrity. The results, denoted as ($\diamond$ w/o Feature F), reveal significant performance declines. Specifically, there is a drop of 3.26% and 2.36% in the applicability and compliance tasks, respectively, for Llama2-13B (see Table 3). These findings demonstrate the importance of feature integrity.
@@
We remove the diversity ranking ($\diamond$ w/o Diversity R) and randomly sample cases for each norm. Low diversity often results in high similarity among cases, such as in the roles of entities or specific categories of information. The lack of diversity can decrease the robustness of training, as demonstrated in Wang et al. (2024b, 2023). This impact is further reflected in a 3.86% decline in the Macro F1-score for applicability judgments in Llama2-13B. Furthermore, we deactivate all of the above filters and ranking mechanisms ($\diamond$ w/o All Parts) and observe significant decreases across all language models, with Mistral-7B experiencing drops of 5.89% and 5.05% in each task, respectively. These findings underscore the importance of enhancing the integrity, consistency, and diversity of generated cases.
@@ 6 Conclusion
Our methodology rigorously adheres to the permit and forbid norms delineated within HIPAA; however, it fails to incorporate the interconnections among these norms. Legal norms frequently entail cross-references, wherein adjudication of cases may hinge upon multiple norms simultaneously, as evidenced by the formalization of legal reasoning Eisenberg (2022); Lam et al. (2009). Future work should construct cases based on multiple norms to reflect real-world scenarios better and potentially yield improvements. Additionally, we do not consider the few-shot setting due to multiple examples often exceeding the maximum input length of LLMs. For the selection of laws, we conduct experiments on HIPAA due to its prominence in the privacy domain and the relatively abundant availability of open-source cases, which can serve as ground truth for testing. We invite legal professionals with access to cases related to other privacy laws to contact us, as this would facilitate the extension of our approach to additional privacy regulations such as COPPA Aftab and Savitt (1999), GDPR Voigt and Von dem Bussche (2017), etc. Moreover, this paper primarily focuses on case generation, and we do not employ techniques such as retrieval-augmented generation Gao et al. (2024); Lewis et al. (2020) or vector embedding Douze et al. (2024) for retrieving relevant norms. We believe that dynamically indexing Liu (2022) and retrieving related norms, based on the statute graph constructed in Section 3.1, is a promising direction.
@@
We use the API provided by the official website of the Code of Federal Regulations to access the HIPAA Privacy Rule. We follow contextual integrity theory Nissenbaum (2004) to generate synthetic cases for constructing GoldCoin-hipaa, and manually remove cases that could potentially leak real-world private information. We follow the official usage and access rules of the Caselaw Access Project [^5] https://case.law/about/#usage-access during downloading relevant cases. Human evaluations and annotations are performed by two legal experts to review the quality of synthetic cases and remove cases that contain explicit content such as gore or violence. Annotations are compensated at 15 USD per hour, above the local minimum wage. To the best of our knowledge, this work complies with open-source agreements and does not pose risks of information leakage or other hazards.
Acknowledgements
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong.